OpenStudy (anonymous):

For the function f(x) = x + 7, what is the ordered pair for the point on the graph when x = 2b A. ( 2b, 2b+7) B. ( 2b, x + 7) C.( x, x + 7 ) D. (x, 2b + 7)

OpenStudy (ajprincess):

in the ordered pair the first value will be the value of x and the second value will be the value of f(x). u can get the value of f(x) by plugging the given value of x in f(x). Does that help?

OpenStudy (anonymous):

Kinda but how do you get the value of x

OpenStudy (ajprincess):

It is given

OpenStudy (anonymous):

so the answer is b.

OpenStudy (ajprincess):

nope. find the value of f(x) when x=2b

OpenStudy (math&ing001):

The pair looks like this (x;y) with y=f(x). Gues again

OpenStudy (math&ing001):

guess*

OpenStudy (anonymous):

i am confused but is it answer d then

OpenStudy (ajprincess):

nope

OpenStudy (ajprincess):

what is the value of x given>

OpenStudy (math&ing001):

You replace x with 2b

OpenStudy (math&ing001):

You replace x with 2b on everything. There shouldn't be any x left in your final answer.

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

can you help me with one more

OpenStudy (math&ing001):

Ok

OpenStudy (math&ing001):

Btw what answer did you get for the first ?

OpenStudy (anonymous):

Find the f(5) f(1) = 2 and f(2) = 4, f(1) + f (2) +f(n-1), for n > 2 so f(5) =

OpenStudy (anonymous):

i got for the first one A.

OpenStudy (math&ing001):

That's correct. Gz!

OpenStudy (math&ing001):

There's something wrong about this part f(1) + f (2) +f(n-1)

OpenStudy (anonymous):

oh yeah before the f(1) there should have been a f(n)= so like 4,f(n) = f(1)

OpenStudy (math&ing001):

You should find f(3) and f(4) before you can find f(5). You just replace n with 3 on this f(n) = f(1) + f (2) +f(n-1) and replace f(1) and f(2) with their respectives values. Same for f(4).

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

for the first one

OpenStudy (anonymous):

well actually its A but thanks for answering

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i knew that

OpenStudy (anonymous):

Then why did you say D.

OpenStudy (anonymous):

by accident

OpenStudy (anonymous):

not kidding

OpenStudy (anonymous):

No Problem :)